OpenStudy (anonymous):

Solve and show the answer in simplest terms: 2/5 and 1/3

OpenStudy (dumbcow):

if we are adding 2/5 + 1/3 = 6/15 + 5/15 = 11/15

OpenStudy (dumbcow):

or 2/5 * 1/3 = 2/15 2/5 / 1/3 = 2/5 * 3/1 = 6/5